PLEASE BRIEF FAZANS RULE/........

Grade:12th Pass

1 Answers

Ashwin Sinha
12 years ago

Dear Amrendra,

inorganic chemistryFajans' Rules, formulated by Kazimierz Fajans in 1923,are used to predict whether a chemical bond will becovalent or ionic, and depend on the charge on the cation and the relative sizes of the cation and anion. They can be summarized in the following table:

IonicCovalent
Low positive charge High positive charge
Large cation Small cation
Small anion Large anion

 

Thus sodium chloride (with a low positive charge (+1), a fairly large cation (~1 Å) and relatively small anion (2Å) is ionic; butaluminium iodide (AlI3) (with a high positive charge (+3) and a large anion) is covalent.

Polarization will be increased by:

  • High charge and small size of the cation
    • Ionic potential Å Z+/r+ (= polarizing power)
  • High charge and large size of the anion
    • The polarizability of an anion is related to the deformability of its electron cloud (i.e. its "softness")
  • An incomplete valence shell electron configuration
    • Noble gas configuration of the cation produces better shielding and less polarizing power
      • e.g. Hg2+ (r+ = 102 pm) is more polarizing than Ca2+ (r+ = 100 pm)

If we first consider an ionic bond, we see that the electron(s) is(are) not shared but transferred between the atoms conveying definite charges to each participant in the bond. Now, we consider the "size" of the charge. I.E. an aluminum atom with a +3 charge has a relatively large positive charge. That positive charge then "tugs" on the electron cloud of the other participant in the bond.

Now, let us consider Aluminum Iodide. I would say that it is an ionic bond with much covalent character. In the AlI3 bonding, the Aluminum gains a +3 charge. The large charge pulls on the electron cloud of the iodines. Now, if we consider the iodine atom, we see that it is relatively large and thus the outer shell electrons are relatively well shielded from the nuclear charge. In this case, the aluminum ion's charge will "tug" on the electron cloud of iodine, drawing it closer to itself. As the electron cloud of the iodine nears the aluminum atom, the negative charge of the electron cloud "cancels" out the positive charge of the aluminum cation. This produces an ionic bond with covalent character.

Now, if we take a different example, for example AlF3. We see that a similar situation occurs, but instead of Iodine we now have fluorine, a relatively small highly electronegative atom. The Fluorine's electron cloud is definitely less shielded from the nuclear charge and will thus be less polarizable. Thus, we get an ionic compound (metal bonded to a nonmetal) with slight covalent character.

 

 

 

Chart illustrating the relationship between atomic and ionic radius
               Good Luck!!!!!!!!
  Plz. approve by clicking 'Yes' given below!!!!!!!!!!   Plz. Plz. Plz.!!!!!!!!!!!